X baskets of strawberries were put on sale on Monday. The unsold baskets of any day were put on sale the next day. 1/3 of the baskets were sold on Monday. 2/3 of those put on sale on Tuesday were unsold. 8 baskets were put on sale on Wednesday. What is X?
Answer by edjones(7311)

2x/3 were put on sale on tues.
2x/3*2/3=4x/9 were put on sale on wed.
4x/9=8
4x=72
x=72/4
x=18
Check:
1/3 of 18 were sold on mon.
12 were put on sale on tues.
2/3 of 12 were unsold so there was 8 left to be put on sale wed.
Ed